# Scalable Hypergraph Structure Learning with Diverse Smoothness Priors

### v1

@@ V-B Lung Cancer Mortality Rates
HSLS, when compared with the baseline K-NN learned hypergraph, always produced smoother structures on the mortality data, even in the presence of more hyperedges. The Sum-Square and Max-Square total variations generate the largest disparity between the baseline and the algorithmic total smoothness measure, which aligns with the discussion of the total variation structural properties in Sec. II. Furthermore, we note that during the learning process, the sets of hyperedges learned by K-NN and HSLS always had overlap. This makes sense as K-NN promotes local smoothness which can occasionally be optimal for global smoothness as well. The total variation using the smoking data is across the board much greater than that of the baseline and algorithmic approaches. However, comparing the metrics of the smoking and temperature data, we find that the correlated smoking signal was always significantly smoother on the original hypergraph than the less correlated temperature. Therefore, we believe that the learned hypergraph could be used to distinguish between potentially correlated and uncorrelated variables to the initial dataset.
V-C Voltage Meter Readings
Thus far, our experiments have demonstrated HSLS to be accurate and globally smooth, specifically when paired with Max-Square variation. From this basis, we now establish our method to also be fast and scalable. Here, a synthetic time-series dataset has been simulated and generated using OpenDSS [34] software on an IEEE34 power circuit, which is a standard test feeder circuit simulating a power distribution system [35]. The dataset is comprised of 95 nodes with over 2000 hourly samples. Each node in the time-series represents a node on the feeder circuit where voltage magnitude has been measured with the help of a smart meter or an Advanced Metering Infrastructure (AMI) [36]. Each of these signals have been assigned a phase A, B, or C with phases differing from each other by 120 degrees. Therefore, this dataset is suitable for clustering problems based upon phase identification where nodes with similar signals are expected to be from the same phase grouping. It is imperative to assign accurate phase labels in order to avoid the situation of an unbalanced three-phase load being delivered to customers.
We experiment on this synthetic data by applying the HSLS with Max-Square variation to demonstrate the scalability and speed with which a hypergraph can be learned on a larger dataset. Previously, we have tested on datasets ranging from 21 to 61 nodes and a maximum cardinality not exceeding $M=3$. Now, we have a much larger 95 node dataset where we learn a uniform hypergraph ranging from $M=3$ to $M=6$. We also set $\hat{\mathcal{K}}=\{10\}$ for Algorithm 1 for each learned hypergraph. Table V reports the time in seconds that Algorithm 1 and Algorithm 2 took to run.
We see that the largest time dependency is on the selection of the valid hyperedge set. While this may seem constraining, due to the nature of the learning problem, Algorithm 1 needs to only run once while the hypergraph topology is expected to remain static. Therefore, after the initial creation of the hyperedge selections, the only time constraint is on Algorithm 2 which, even in high cardinality scenarios, remains under 750ms runtime. Determining a fast speed is certainly application dependent, however, in this particular scenario, these times are acceptable as the observations are taken hourly, meaning our method has learned the hypergraph well before the next set of observations are completed.
[…]
TABLE V: Time results of uniform hypergraph learning using voltage metering.
$M=3$ $M=4$ $M=5$ $M=6$
Algorithm 1 Time (s) 8.62 56.19 161.21 220.98
Algorithm 2 Time (s) 0.15 0.44 0.68 0.74
@@ VI Conclusion
In this paper, we developed a hypergraph learning optimization method, HSLS, where we aimed to overcome shortcomings of previous works including the lack of a smoothness prior, broad usage of total variation, and algorithmic assumptions. We overcame the first two problems by basing our optimization framework on minimizing the total variation while allowing the smoothness term to be robust for substitution. The choice to format our optimization with a primal and dual variable and the investigation into the use of K-NN as a means of hyperedge possibility reduction solved the issues surrounding the algorithmic assumptions related to weighted hyperedge importance and valid hyperedge search spaces. Furthermore, the K-NN reduction method we improved upon made our method scalable without compromising the span of the hyperedge search. We were able to experimentally evaluate the effectiveness of our proposed method, with each selected total variation, in recovering a hypergraph structure based on the smoothness prior.
When investigating our method’s ability to accurately recover a hypergraph compared against the ground truth topology, we found that our proposed method HSLS with the Max-Square was the overall best performer when compared to the other smoothness terms and the results of other methods. This validated our optimization framework showing that reasonable hypergraph structure recovery was possible and total variation substitution was viable. We proceeded with evaluating how well our method performed on encouraging a smooth hypergraph structure. We found that across all total variations, our method encouraged global smoothness more so than the K-NN baseline. We also determined that our method recovers a structure that is smoother across data correlated to the original dataset and less smooth across uncorrelated data, highlighting a key property of the learned hypergraph structure. Finally, we demonstrated our method’s ability to scale to larger problems with more nodes and significantly larger maximum cardinalities. We also determined that the main time dependence was on the hyperedge reduction algorithm, but this was of little issue as the expectation is to complete that process only once per static application.
In future work, we will extend concepts to a dynamic domain where time-series data is subject to alter as more observations are taken and where the hypergraph structure at one point in time may not be consistent with future points in time. We believe that our work here is a suitable starting point to evolve the research in this direction.

### v2

@@ V-B Lung Cancer Mortality Rates
HSLS, when compared with the baseline K-NN learned hypergraph, always produced smoother structures on the mortality data, even in the presence of more hyperedges. The Sum-Square and Max-Square total variations generate the largest disparity between the baseline and the algorithmic smoothness measure. Furthermore, we note that during the learning process, the sets of hyperedges learned by K-NN and HSLS always had overlap. This makes sense as K-NN promotes local smoothness which can occasionally be optimal for global smoothness as well. The total variation using the smoking data is across the board much greater than that of the baseline and algorithmic approaches. However, comparing the metrics of the smoking and temperature data, we find that the correlated smoking signal was always significantly smoother on the original hypergraph than the less correlated temperature. Therefore, we believe that the learned hypergraph could be used to distinguish between potentially correlated and uncorrelated variables to the initial dataset.
V-C Hyperspectral Image Classification
A hyperspectral image (HSI) captures spectral information across a range of wavelengths, including those outside of the visible light spectrum. We selected a HSI benchmark dataset known as Indian Pines to perform classification on. This data was originally captured by the Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) sensor over the Purdue University Agronomy farm [53], though for this work we use the filtered version of the HSI where 20 bands related to water absorption are removed [54]. This leaves us with a HSI of 145 x 145 pixels, each with 200 spectral reflectance band features in the 400-2500 nm wavelength range. Figure 7a contains a false-color depiction of the HSI being tested on. Each pixel has been assigned a class label 0-16 where label 0 is a null class and will be omitted from here on. The remaining classes 1-16 can be identified by name and quantity in Table IV. Cross referencing the colors of Table IV with Fig. 7b reveals the location of pixel classes in the original HSI.
Figure 7: Images based on HSI data: (a) False-color pixel map, (b) Ground truth pixel map with black color for null pixels.
The goal of this experiment is to use HSLS to improve upon the results of an existing hypergraph based neural network to demonstrate the method’s applicability and continue observing the impact of the diverse total variation measurements. The existing neural network framework we chose from Ma et al. [55], known as $\text{F}^{2}\text{HNN}_{SCC}$, is based on fusion of spectral, spatial, and convolutional neural network (CNN) features. The general framework starts with pre-training a 3-D CNN to generate $\mathbf{X}_{CNN}$ with features for each pixel. Then, the spatial coordinate location of the pixels are used to populate $\mathbf{X}_{spa}$, and the flattened features of the HSI define $\mathbf{X}_{spe}$ the spectral features. From these three generated features sets, corresponding hypergraph incidence matrices $\mathbf{H}_{spe},\mathbf{H}_{spa}$, and $\mathbf{H}_{CNN}$ are constructed from a radial basis function. The fusion occurs by concatenating $\mathbf{H}=[\mathbf{H}_{spe},\mathbf{H}_{spa},\mathbf{H}_{CNN}]$ where $\mathbf{H}$ inputs into a hypergraph convolution layer as
$\mathbf{X}^{s+1}=\sigma(\mathbf{D}_{v}^{-1/2}\mathbf{H}\mathbf{W}\mathbf{D}_{e}^{-1}\mathbf{H}^{\top}\mathbf{D}_{v}^{-1/2}\mathbf{X}^{(s)}\mathbf{\Theta}^{(s)})$ (23)
where $\mathbf{X}^{s+1}$ is the output of the $s$-th layer, $\sigma$ is the nonlinear activation, $\mathbf{D}_{v}$ is the diagonal node degree matrix with diagonal elements $d_{v_{i}}=\sum_{\mathbf{e}_{j}\in E(\mathcal{H})}w_{\mathbf{e}_{j}}\mathbf{H}_{ij}$, $\mathbf{D}_{e}$ is the diagonal hyperedge degree matrix with diagonal elements $d_{e_{j}}=\sum_{v_{i}\in V(\mathcal{H})}\mathbf{H}_{ij}$, and $\mathbf{W}$ is the diagonal hyperedge weight matrix with diagonal elements $\mathbf{w}$. $\text{F}^{2}\text{HNN}_{SCC}$ passes $\mathbf{H}$ through two hypergraph convolution layers, where softmax is then applied row-wise to get the final classification predictions. We refer the readers to [55] for a complete view of the neural network architecture as we implement their work directly.
TABLE IV: Indian Pines Data Splits.
Class Class Color Class Name Training Testing
1   Alfalfa 15 31
2   Corn Notill 50 1378
3   Corn Mintill 50 780
4   Corn 50 187
5   Grass Pasture 50 433
6   Grass Trees 50 680
7   Grass Pasture Mowed 15 13
8   Hay Windrowed 50 428
9   Oats 15 5
10   Soybean Notill 50 922
11   Soybean Mintill 50 2405
12   Soybean Clean 50 543
13   Wheat 50 155
14   Woods 50 1215
15   Buildings Grass Trees Drives 50 336
16   Stone Steel Towers 50 43
Total 695 9554
We note that superpixel segmentation using Simple Linear Iterative Clustering (SLIC) was added as a pre-processing step, which is commonly used in such applications [56, 57]. The superpixels are comprised of groups of pixels segmented by SLIC and thereby reduce the spatial dimension. So, for our implementation, the 145 x 145 pixel map was reduced to 194 superpixels. We also update $\mathbf{X}_{spe}\in\mathbb{R}^{194\times 200}$ and $\mathbf{X}_{CNN}\in\mathbb{R}^{194\times 256}$ where each row is the average sum of the features from the pixels in each superpixel group. $\mathbf{X}_{spa}\in\mathbb{R}^{194\times 2}$ now contains the centroid locations of the superpixels.
TABLE V: Indian Pines classification results on baseline method $\text{F}^{2}\text{HNN}_{SCC}$ and several implementations of HSLS.
Class $\textbf{F}^{2}\textbf{HNN}_{SCC}$ [55] Uni Sum-Square Uni Sum-Abs Uni Max-Abs Uni Max-Square Non-Uni Sum-Square Non-Uni Sum-Abs Non-Uni Max-Abs Non-Uni Max-Square
1 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44 97.42 $\pm$ 1.44
2 69.07 $\pm$ 9.12 83.88 $\pm$ 8.15 82.31 $\pm$ 10.70 83.35 $\pm$ 6.36 86.91 $\pm$ 5.24 81.28 $\pm$ 9.31 85.54 $\pm$ 5.51 80.46 $\pm$ 5.20 84.80 $\pm$ 8.04
3 77.82 $\pm$ 26.19 88.08 $\pm$ 5.53 78.13 $\pm$ 14.26 81.44 $\pm$ 8.41 86.49 $\pm$ 7.79 88.54 $\pm$ 5.68 86.56 $\pm$ 7.69 62.33 $\pm$ 10.33 88.95 $\pm$ 3.49
4 94.33 $\pm$ 2.66 93.69 $\pm$ 1.33 93.05 $\pm$ 0.76 93.05 $\pm$ 0.76 93.05 $\pm$ 0.76 94.01 $\pm$ 1.62 93.37 $\pm$ 1.39 96.26 $\pm$ 0.76 93.69 $\pm$ 1.33
5 69.61 $\pm$ 24.01 87.76 $\pm$ 5.97 86.56 $\pm$ 4.19 88.73 $\pm$ 5.30 92.06 $\pm$ 4.56 88.82 $\pm$ 4.52 88.73 $\pm$ 6.19 89.61 $\pm$ 6.02 87.67 $\pm$ 6.83
6 93.56 $\pm$ 4.15 89.15 $\pm$ 4.06 89.09 $\pm$ 4.04 89.21 $\pm$ 4.10 89.09 $\pm$ 4.04 89.03 $\pm$ 4.00 89.03 $\pm$ 4.00 84.35 $\pm$ 7.56 89.03 $\pm$ 4.00
7 92.31 $\pm$ 17.20 84.62 $\pm$ 21.07 92.31 $\pm$ 17.20 78.46 $\pm$ 19.91 70.77 $\pm$ 16.68 84.62 $\pm$ 21.07 76.92 $\pm$ 21.07 66.15 $\pm$ 40.92 84.62 $\pm$ 21.07
8 99.21 $\pm$ 1.09 99.16 $\pm$ 1.15 99.21 $\pm$ 1.09 99.58 $\pm$ 0.94 99.16 $\pm$ 1.15 99.58 $\pm$ 0.94 99.58 $\pm$ 0.94 98.36 $\pm$ 0.92 99.16 $\pm$ 1.15
9 60.00 $\pm$ 54.77 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 72.00 $\pm$ 38.99 100.00 $\pm$ 0.00
10 68.16 $\pm$ 25.29 78.31 $\pm$ 5.20 79.48 $\pm$ 9.66 79.57 $\pm$ 4.72 75.29 $\pm$ 8.88 76.62 $\pm$ 11.24 79.11 $\pm$ 1.72 80.98 $\pm$ 3.27 76.36 $\pm$ 8.77
11 68.30 $\pm$ 20.79 87.44 $\pm$ 5.62 90.38 $\pm$ 5.15 88.42 $\pm$ 6.54 87.64 $\pm$ 4.02 88.62 $\pm$ 3.87 90.41 $\pm$ 3.92 78.55 $\pm$ 14.12 87.81 $\pm$6.18
12 78.97 $\pm$ 22.88 86.70 $\pm$ 6.14 88.77 $\pm$ 3.89 88.77 $\pm$ 3.89 87.74 $\pm$ 3.45 83.20 $\pm$ 7.58 87.04 $\pm$ 6.28 76.21 $\pm$ 3.13 86.23 $\pm$ 5.79
13 98.71 $\pm$ 0.91 98.71 $\pm$ 0.91 98.71 $\pm$ 0.91 98.71 $\pm$ 0.91 98.71 $\pm$ 0.91 98.71 $\pm$ 0.91 98.71 $\pm$ 0.91 99.23 $\pm$ 0.71 98.71 $\pm$ 0.91
14 97.70 $\pm$ 2.91 98.42 $\pm$ 2.49 99.56 $\pm$ 0.99 99.56 $\pm$ 0.99 98.42 $\pm$ 2.49 99.54 $\pm$ 0.99 99.54 $\pm$ 0.99 94.03 $\pm$ 0.99 99.54 $\pm$ 0.99
15 98.04 $\pm$ 0.27 98.75 $\pm$ 1.14 98.45 $\pm$ 0.90 98.04 $\pm$ 0.27 98.75 $\pm$ 1.14 93.81 $\pm$ 9.52 98.45 $\pm$ 0.90 98.33 $\pm$ 0.93 99.17 $\pm$ 1.14
16 79.07 $\pm$ 44.22 98.61 $\pm$ 1.27 98.61 $\pm$ 1.27 98.61 $\pm$ 1.27 98.61 $\pm$ 1.27 98.61 $\pm$ 1.27 98.61 $\pm$ 1.27 98.61 $\pm$ 1.27 98.61 $\pm$ 1.27
OA (%) 78.98 $\pm$ 8.97 88.90 $\pm$ 0.87 88.91 $\pm$ 1.72 88.93 $\pm$ 1.17 89.19 $\pm$ 0.69 88.53 $\pm$ 3.06 90.03 $\pm$ 0.36 82.88 $\pm$ 3.95 89.13 $\pm$ 0.82
AA (%) 83.89 $\pm$ 10.85 91.92 $\pm$ 1.01 92.00 $\pm$ 1.24 91.43 $\pm$ 1.22 91.26 $\pm$ 1.43 91.40 $\pm$ 1.40 91.81 $\pm$ 1.06 85.81 $\pm$ 2.27 91.98 $\pm$ 1.21
$\kappa$ 0.762 $\pm$ 0.10 0.873 $\pm$ 0.01 0.873 $\pm$ 0.02 0.874 $\pm$ 0.01 0.876 $\pm$ 0.01 0.869 $\pm$ 0.04 0.886 $\pm$ 0.00 0.806 $\pm$ 0.04 0.876 $\pm$ 0.01
We implement $\text{F}^{2}\text{HNN}_{SCC}$ as the baseline method with the recommended radial basis function to generate weighted $\mathbf{H}$. This is compared against the use of HSLS to generate binary $\mathbf{H}$ with $\mathbf{X}=[\mathbf{X}_{spe},\mathbf{X}_{spa},\mathbf{X}_{CNN}]$ being the signal input for both methods. For HSLS, we recover both a uniform hypergraph structure with $\mathcal{K}=\{3\}$ and a non-uniform hypergraph structure with $\mathcal{K}=\{2,3,4\}$. All proposed total variation terms are tested to recover $\mathbf{w}$, which in turn was used to populate the binary $\mathbf{H}$, and these terms together can generate $\mathbf{D}_{v}$, $\mathbf{D}_{e}$, and $\mathbf{W}$. Algorithm 1 was used on $\mathbf{X}$ to reduce the hyperedge search space, creating a scalable problem and enabling the feasibility of the non-uniform testing. After generating $\mathbf{H}$, the hypergraph convolutional neural network was applied which outputs classification predictions.
Both the baseline and HSLS methods have the same training and testing scheme and were evaluated on the same training and testing splits. We note that the selected training method is derivative of other works in HSI classification [55, 58]. Table IV reports the number of training and testing samples per class. We implemented a Monte Carlo cross-validation setup such that, for 5 total passes, the training and testing splits were randomized, but adhered to Table IV. Then, 4 validation samples per class were randomly sampled from the training data. The neural network model was trained over 300 epochs using the Adam optimizer and cross entropy loss, starting at a learning rate of 0.01 and decreasing based on a learning rate scheduler. Every epoch, the model was evaluated on the validation set and the validation loss was recorded. For more details, we refer the reader to the code provided. After training, the model with the minimum validation loss was used for testing. The test results were averaged across 5 trials and are reported in Table V where the baseline is $\text{F}^{2}\text{HNN}_{SCC}$, and all other columns are HSLS either uniform (written as Uni) or non-uniform (written as Non-Uni) combined with one of the four total variation terms (with Absolute written as Abs). The majority of the table reports accuracy per class, OA is the overall accuracy, AA is the average accuracy found from the mean of the sum of individual class accuracies, and $\kappa$ is Cohen’s kappa value, a measurement of how closely the model agrees with the ground truth labels while accounting for random chance.
The use of HSLS across all variants always improves upon the OA, AA, and $\kappa$ of the baseline value from $\text{F}^{2}\text{HNN}_{SCC}$. Furthermore, the per class accuracy is also generally improved, with few exceptions. This justifies the use of HSLS as a viable alternative for hypergraph structure inference for use in downstream tasks. Comparing the results among the HSLS models, the best performer could be attributed to the non-uniform Sum-Absolute implementation, though many of the overall results are within a percent or two of each other, with the worst result coming from the non-uniform Max-Absolute model. This is a striking difference to the results of the co-authorship networks, as all total variation terms seem to produce viable outputs for classification. This could be contributed to the fluidity of the HSI structure where, in the absence of a ground truth hypergraph, there are many hypergraphs that could appropriately describe some underlying topology. This supports the conclusion that the effectiveness of hypergraph total variation terms are application dependent.
@@ VI Conclusion
In this paper, we developed a novel interpretation of a hypergraph learning optimization method, HSLS, which aimed to remedy the lack of a smoothness prior, shortage of total variation testing, algorithmic assumptions, and scalability. We overcame the first two problems by basing our optimization framework on minimizing the total variation while allowing the smoothness term to be robust for substitution. A diverse set of smoothness terms were introduced such that the efficacy of a range of definitions could be tested in different applications. The choice to format our optimization with a primal and dual variable and the investigation into the use of K-NN as a means of hyperedge possibility reduction solved the issues surrounding the algorithmic assumptions related to weighted hyperedge importance and valid hyperedge search spaces. Furthermore, the K-NN reduction method we improved upon made our method scalable without compromising the span of the hyperedge search. We were able to experimentally evaluate the effectiveness of our proposed method, with each selected total variation, in recovering a hypergraph structure and thereby generated original empirical results.
When investigating our method’s ability to accurately recover a hypergraph with reference to the ground truth topology, we found that our proposed method HSLS with Max-Square was the overall best performer when compared to the other smoothness terms and the results of other methods. This validated our optimization framework showing that reasonable hypergraph structure recovery was possible and total variation substitution was viable. Furthermore, this supported the conclusion that Max-Square, of the tested total variations, would be the best choice of smoothness in the application of recovering a specific hypergraph structure.
We proceeded with evaluating how well our method encouraged smooth hypergraph structures. We found that, across all total variations, our method encouraged global smoothness more so than the K-NN baseline. We also determined that our method recovers a structure that is smoother across data correlated to the original dataset and less smooth across uncorrelated data, highlighting a key property of the learned hypergraph structure.
Finally, we demonstrated our method’s ability to scale to larger applications, defined by more nodes and higher cardinalities, through improving upon an existing hypergraph based method of hyperspectral image classification. The metric report demonstrated the performance increase of using HSLS over the baseline and revealed that, in a domain where the ground truth hypergraph is not present, all total variation terms across uniform and non-uniform scenarios were viable candidates for better results, which is similar to the outcome of the mortality rate experimentation.
In future work, we will extend concepts to a dynamic domain where time-series data is subject to alter as more observations are taken. This implies that the hypergraph structure at one point in time may not be consistent with future points in time. We believe that our work here is a suitable starting point to evolve the research in this direction.